# From Word Vectors to Multimodal Embeddings: Techniques, Applications, and Future Directions For Large Language Models

### v1

@@ III-C2 Challenges and Evaluation of Cross-Lingual Embeddings
Developing cross-lingual embeddings entails significant challenges, especially in low-resource settings and when data domains differ across languages. Limited training data can impair unsupervised methods, as noted in [60], where application to the language Yongning Na highlights resource scarcity challenges. Domain differences across languages also impact embedding quality; while [60] demonstrates that their model can handle different-domain corpora, achieving high-quality embeddings remains challenging.
Evaluating cross-lingual embeddings typically involves both intrinsic and extrinsic tasks. Intrinsic evaluations assess embeddings through word similarity tasks, while extrinsic evaluations examine embedding performance in downstream applications. For instance, [30] uses the NIST08 Chinese-English machine translation task for extrinsic evaluation. For a thorough discussion on the challenges and evaluation of cross-lingual embeddings, [61] provides extensive analysis.
@@ III-C3 Language-Agnostic Sentence Embeddings: Techniques, Evaluation, and Addressing Bias
Language-agnostic sentence embeddings represent sentences from different languages in a unified vector space, enabling cross-lingual semantic similarity and retrieval tasks. Techniques for generating these embeddings include knowledge distillation and fine-tuning of pre-trained multilingual models. [52] proposes extending monolingual sentence embeddings to new languages via knowledge distillation, where a student model replicates the vector space of a monolingual teacher model. This approach shows strong results for low-resource languages and effectively minimizes language bias. Additionally, [29] introduces LaBSE, a language-agnostic BERT sentence embedding model that achieves state-of-the-art performance on bi-text retrieval tasks.
Evaluation of language-agnostic sentence embeddings often involves tasks like semantic textual similarity (STS) and bi-text mining. Addressing language bias, where models may preferentially align certain language pairs, is a key challenge. [52] notes that their approach, as well as mUSE, avoid language bias issues seen in models like LASER and LaBSE.
@@ IV-A2 Methods for Visual Grounding
A common approach to visual grounding involves creating joint embedding spaces for images and text. In these spaces, visual and textual data are mapped into a shared vector representation, enabling direct comparison and interaction between the two modalities. [63] introduces an encoder-decoder pipeline to achieve this. In this architecture, an encoder maps both images and sentences into a shared latent space, and a decoder generates descriptions from this space. Another approach utilizes multimodal contrastive learning, as described in [64], where a two-stream model for visual and language inputs aligns representations by maximizing agreement between paired image-text data while minimizing agreement for unrelated pairs.
@@
Attention mechanisms are highly effective for aligning visual and textual information in cross-modal retrieval tasks. By employing attention, models can identify the most relevant visual features for a textual query or vice versa. [64] employs a cross-modal attention module that retrieves specific visual objects based on language queries. This mechanism allows for fine-grained alignment, enabling the model to focus on essential parts of the input specific to a task.
@@
Similarity learning plays an essential role in grounding representations of objects, verbs, and attributes. [65] uses similarity learning within an embodied simulation environment, enabling comparison of objects based on their properties and behaviors. By learning a similarity metric through interactions within this simulation, the model effectively grounds word vectors to object representations, facilitating tasks like zero-shot learning, where novel objects can be recognized based on descriptions alone. Notably, [65] also found that grounding object representations enhances the grounding of verbs and attributes, aligning with findings in analogical reasoning.
@@ IV-A3 Applications of Visually Grounded Language Models
Image captioning generates textual descriptions for images, while image retrieval involves finding images that match a given textual query. [63] introduced an encoder-decoder pipeline that creates a joint embedding space for images and text, combined with a ”structure-content” neural language model. This model achieved state-of-the-art results in both image captioning and retrieval tasks on the Flickr8K and Flickr30K datasets.
@@
Visual Question Answering (VQA) requires answering questions based on visual information within images. [35] explored multimodal embeddings by integrating language models with acoustic information and paralinguistic cues to enhance emotion recognition—a crucial aspect for VQA applications. Additionally, [66] developed a deep neural network that learns a shared embedding space for point-cloud, natural language, and manipulation trajectory data, enabling robots to manipulate novel objects based on instructions and visual experience. [67] introduced the Visual Semantic Embedding Probe (VSEP) to evaluate contextualized word embeddings for zero-shot object recognition in complex visual scenes. Further, [68] utilized pre-trained language models to create interpretable and efficient predictors for tasks, including visually grounded applications.
@@
Referring expression comprehension identifies an object in an image based on a textual description, while referring expression generation produces a unique textual description for a specified object within an image. [69] introduced Kosmos-2, a multimodal large language model capable of grounding text to visual information, such as drawing bounding boxes around objects. Trained on a large-scale dataset of grounded image-text pairs, Kosmos-2 demonstrated enhanced performance on tasks involving referring expression comprehension and generation.
@@
Multimodal image search uses a blend of textual and visual queries to locate images. [64] proposed a two-stream model that aligns visual and language representations through cross-modal contrastive learning, enabling multimodal image searches using queries composed of images, text, or combinations of both.
@@ IV-A4 Impacts and Challenges
Visually grounded language models have demonstrated benefits across a range of linguistic tasks. [70] investigated the impact of visual grounding on both concrete and abstract words, finding that grounding can enhance representations for both types of words. This study also highlighted the advantages of visual grounding for contextualized embeddings, particularly when trained on smaller corpora, where the added visual context helps compensate for limited textual data. [15] examined the role of contextual information in language modeling, showing that context is essential for addressing semantic challenges in tasks like word sense disambiguation.
Despite these advancements, visually grounded language models face notable challenges. [67] identified the lack of mutual exclusivity bias as a limitation in current visual semantic embedding models, where models may fail to distinguish between overlapping or closely related objects within scenes. Addressing such biases and enabling language models to reason more effectively about novel visual scenes are key areas for future research, with the goal of enhancing the models’ robustness and generalization across diverse visual-linguistic contexts.
@@ IV-B1 Motivation: Challenges in Representing and Integrating Diverse Sensor Modalities
Robots operating in unstructured, real-world environments face significant challenges in processing and integrating data from diverse sensor modalities. These include visual data from cameras and depth sensors, natural language instructions from human operators, haptic feedback, and motion trajectories from manipulators. Integrating these varied data streams is essential for robots to build a comprehensive understanding of their environment and interact effectively. However, as highlighted by [66], manually designing features that bridge these modalities is extremely difficult due to differences in data formats and inherent complexities. For instance, visual data may be represented as images, point clouds, or voxel grids, each with unique dimensional structures, whereas language instructions are symbolic sequences that must be translated into a format suitable for robotic control. Similarly, manipulation trajectories can be represented as time series data of joint angles, end-effector positions, or force/torque measurements. This heterogeneity presents a substantial barrier to direct comparison or fusion of information across modalities.
Developing techniques that can automatically learn joint representations to capture relationships between these diverse modalities is crucial for advancing robot autonomy and intelligence. Additionally, as noted by [15], context is essential for interpreting both linguistic and sensorimotor information. While embeddings are effective in capturing syntactic and some semantic relationships, they require contextualization to handle challenges such as word sense disambiguation and the integration of real-world knowledge, particularly in robotics applications.
@@ IV-B2 Methods for Multimodal Embeddings in Robotics
Deep neural networks are a powerful tool for learning shared embedding spaces that unify different modalities, mapping them into a common vector space where semantically related concepts are closer together. Various architectures are used for this purpose. For example, [64] discusses a two-stream architecture where separate networks process individual modalities, such as vision and language, before merging outputs into a joint representational space. Alternatively, a single network can concurrently process all modalities, directly learning a shared embedding space.
Training strategies for multimodal embeddings typically involve a two-stage process: an initial pre-training phase on large unimodal datasets, such as image-caption pairs or text corpora, to learn general features within each modality, followed by fine-tuning on smaller multimodal datasets to align representations across modalities. This approach proved effective in [35], where acoustic information was integrated into contextualized lexical embeddings to enhance emotion recognition through paralinguistic cues.
@@
Effective multimodal embedding training relies on carefully designed loss functions. Margin-based approaches, which aim to separate embeddings of dissimilar concepts while grouping similar ones, are commonly used. For instance, triplet loss, as discussed by [70], teaches the model to position an anchor instance closer to a positive (semantically similar) instance than to a negative (semantically dissimilar) instance by a predefined margin. Similarly, contrastive loss, employed in [66], works by distinguishing between similar and dissimilar pairs, pushing dissimilar embeddings apart while pulling similar ones closer. These loss functions encourage a semantically structured embedding space that effectively captures relationships between modalities.
@@ IV-B3 Combining Point-Cloud Data, Natural Language Instructions, and Manipulation Trajectories
Embedding these diverse modalities into a shared space, as explored in [66], enables the establishment of meaningful connections across representations. For instance, robots can ground language instructions, such as ”grasp the blue object,” by linking the instruction to the corresponding point cloud of a blue object and its associated manipulation trajectory. This multimodal grounding enables generalization to novel objects and instructions, fostering more flexible and adaptive robotic behaviors.
@@ IV-B4 Applications of Multimodal Embeddings in Robotics
Multimodal embeddings enable robots to interact with and manipulate unfamiliar objects based on previous experiences. By learning a shared embedding space that includes object representations (e.g., point clouds), language instructions, and manipulation trajectories, as demonstrated by [66], robots can generalize their knowledge to novel objects. When encountering a new object alongside a language instruction, the robot can map the object’s representation into the shared embedding space, retrieving the most relevant manipulation trajectory aligned with the instruction. This capability allows robots to adapt to new situations and perform tasks without requiring explicit programming for every new object-action combination.
@@
Comprehending and executing complex tasks specified via natural language is a critical objective in robotics. Multimodal embeddings facilitate this by aligning language commands with corresponding action sequences. This enables robots to understand and perform multi-step tasks expressed in natural language, such as ”navigate to the table, pick up the book, and place it on the shelf.” With a multimodal embedding model, the robot can translate these instructions into a sequence of actions, including navigation, object recognition, grasping, and placement. This approach moves beyond simple, keyword-based command interpretation, fostering more flexible and sophisticated human-robot interactions. Future research, as suggested by [40], could enhance this capability by fine-tuning language models with attention mechanisms to improve the accuracy and robustness of mapping natural language instructions to specific actions.
@@
Evaluating multimodal embeddings in robotics requires a mix of real-world experiments and simulations. Real-world experiments provide a realistic measure of a model’s ability to generalize to new scenarios and handle the complexities of unstructured environments. However, these tests can be resource-intensive and time-consuming. Simulated environments offer a controlled and efficient platform to evaluate model performance across various tasks and conditions. For instance, [66] demonstrates how simulations can pre-train models before deploying them on physical robots. Common evaluation metrics include task success rate, execution time, and path efficiency.
@@
The field of multimodal embeddings for robotics presents ongoing challenges and exciting research directions. One primary challenge is scalability to high-dimensional sensor data. As robots incorporate advanced sensors like high-resolution cameras and lidar, the dimensionality of data increases significantly. Efficiently processing and embedding this high-dimensional data is essential for real-world applications. Another critical challenge is achieving real-time performance; real-time sensor data processing and action generation are vital for effective interaction in dynamic environments. Continued research is needed to create more efficient embedding models and architectures that support real-time robotic control. Exploring alternative embedding techniques, such as those discussed by [10], could yield scalable, efficient solutions for handling high-dimensional data in real-time applications.
@@ IV-C1 Motivation: Connecting Artificial and Biological Language Processing
The impressive capabilities of large language models (LLMs) in natural language processing tasks have prompted research into their internal representations and whether these mirror the way the human brain processes language. [32] discusses the shift from word-level to sentence-level representations in language models, which reflects advances in understanding complex language representations. Exploring the connection between artificial and biological language processing is essential for developing AI that is more human-like and for gaining insights into the cognitive foundations of language comprehension. [71] investigates how the structure of language representations correlates with human brain responses, highlighting the potential for identifying shared computational principles. Additionally, [15] emphasizes the role of embeddings and contextual information, which could enhance our understanding of both artificial and biological language processing systems. This line of research not only enriches cognitive neuroscience but also informs the development of more robust and interpretable language models.

### v2

@@ III-C2 Challenges and Evaluation of Cross-Lingual Embeddings
Developing cross-lingual embeddings entails significant challenges, especially in low-resource settings and when data domains differ across languages. Limited training data can impair unsupervised methods, as noted in [takashi_unsupervised_2018], where application to the language Yongning Na highlights resource scarcity challenges. Domain differences across languages also impact embedding quality; while [takashi_unsupervised_2018] demonstrates that their model can handle different-domain corpora, achieving high-quality embeddings remains challenging.
Evaluating cross-lingual embeddings typically involves both intrinsic and extrinsic tasks. Intrinsic evaluations assess embeddings through word similarity tasks, while extrinsic evaluations examine embedding performance in downstream applications. For instance, [will_bilingual_2013] uses the NIST08 Chinese-English machine translation task for extrinsic evaluation. For a thorough discussion on the challenges and evaluation of cross-lingual embeddings, [sebastian_survey_2017] provides extensive analysis.
@@ III-C3 Language-Agnostic Sentence Embeddings: Techniques, Evaluation, and Addressing Bias
Language-agnostic sentence embeddings represent sentences from different languages in a unified vector space, enabling cross-lingual semantic similarity and retrieval tasks. Techniques for generating these embeddings include knowledge distillation and fine-tuning of pre-trained multilingual models. [nils_making_2020] proposes extending monolingual sentence embeddings to new languages via knowledge distillation, where a student model replicates the vector space of a monolingual teacher model. This approach shows strong results for low-resource languages and effectively minimizes language bias. Additionally, [fangxiaoyu_language-agnostic_2020] introduces LaBSE, a language-agnostic BERT sentence embedding model that achieves state-of-the-art performance on bi-text retrieval tasks.
Evaluation of language-agnostic sentence embeddings often involves tasks like semantic textual similarity (STS) and bi-text mining. Addressing language bias, where models may preferentially align certain language pairs, is a key challenge. [nils_making_2020] notes that their approach, as well as mUSE, avoid language bias issues seen in models like LASER and LaBSE.
@@ IV-A2 Methods for Visual Grounding
A common approach to visual grounding involves creating joint embedding spaces for images and text. In these spaces, visual and textual data are mapped into a shared vector representation, enabling direct comparison and interaction between the two modalities. [ryan_unifying_2014] introduces an encoder-decoder pipeline to achieve this. In this architecture, an encoder maps both images and sentences into a shared latent space, and a decoder generates descriptions from this space. Another approach utilizes multimodal contrastive learning, as described in [yizhen_explainable_2021], where a two-stream model for visual and language inputs aligns representations by maximizing agreement between paired image-text data while minimizing agreement for unrelated pairs.
@@
Attention mechanisms are highly effective for aligning visual and textual information in cross-modal retrieval tasks. By employing attention, models can identify the most relevant visual features for a textual query or vice versa. [yizhen_explainable_2021] employs a cross-modal attention module that retrieves specific visual objects based on language queries. This mechanism allows for fine-grained alignment, enabling the model to focus on essential parts of the input specific to a task.
@@
Similarity learning plays an essential role in grounding representations of objects, verbs, and attributes. [ghaffari_grounding_2023] uses similarity learning within an embodied simulation environment, enabling comparison of objects based on their properties and behaviors. By learning a similarity metric through interactions within this simulation, the model effectively grounds word vectors to object representations, facilitating tasks like zero-shot learning, where novel objects can be recognized based on descriptions alone. Notably, [ghaffari_grounding_2023] also found that grounding object representations enhances the grounding of verbs and attributes, aligning with findings in analogical reasoning.
@@ IV-A3 Applications of Visually Grounded Language Models
Image captioning generates textual descriptions for images, while image retrieval involves finding images that match a given textual query. [ryan_unifying_2014] introduced an encoder-decoder pipeline that creates a joint embedding space for images and text, combined with a ”structure-content” neural language model. This model achieved state-of-the-art results in both image captioning and retrieval tasks on the Flickr8K and Flickr30K datasets.
@@
Visual Question Answering (VQA) requires answering questions based on visual information within images. [shao-yen_multimodal_2019] explored multimodal embeddings by integrating language models with acoustic information and paralinguistic cues to enhance emotion recognition—a crucial aspect for VQA applications. Additionally, [jaeyong_deep_2015] developed a deep neural network that learns a shared embedding space for point-cloud, natural language, and manipulation trajectory data, enabling robots to manipulate novel objects based on instructions and visual experience. [yue_language_2021] introduced the Visual Semantic Embedding Probe (VSEP) to evaluate contextualized word embeddings for zero-shot object recognition in complex visual scenes. Further, [singh2023emb] utilized pre-trained language models to create interpretable and efficient predictors for tasks, including visually grounded applications.
@@
Referring expression comprehension identifies an object in an image based on a textual description, while referring expression generation produces a unique textual description for a specified object within an image. [peng_kosmos-2_2023] introduced Kosmos-2, a multimodal large language model capable of grounding text to visual information, such as drawing bounding boxes around objects. Trained on a large-scale dataset of grounded image-text pairs, Kosmos-2 demonstrated enhanced performance on tasks involving referring expression comprehension and generation.
@@
Multimodal image search uses a blend of textual and visual queries to locate images. [yizhen_explainable_2021] proposed a two-stream model that aligns visual and language representations through cross-modal contrastive learning, enabling multimodal image searches using queries composed of images, text, or combinations of both.
@@ IV-A4 Impacts and Challenges
Visually grounded language models have demonstrated benefits across a range of linguistic tasks. [shahmohammadi_language_2022] investigated the impact of visual grounding on both concrete and abstract words, finding that grounding can enhance representations for both types of words. This study also highlighted the advantages of visual grounding for contextualized embeddings, particularly when trained on smaller corpora, where the added visual context helps compensate for limited textual data. [matthias_context_2018] examined the role of contextual information in language modeling, showing that context is essential for addressing semantic challenges in tasks like word sense disambiguation.
Despite these advancements, visually grounded language models face notable challenges. [yue_language_2021] identified the lack of mutual exclusivity bias as a limitation in current visual semantic embedding models, where models may fail to distinguish between overlapping or closely related objects within scenes. Addressing such biases and enabling language models to reason more effectively about novel visual scenes are key areas for future research, with the goal of enhancing the models’ robustness and generalization across diverse visual-linguistic contexts.
@@ IV-B1 Motivation: Challenges in Representing and Integrating Diverse Sensor Modalities
Robots operating in unstructured, real-world environments face significant challenges in processing and integrating data from diverse sensor modalities. These include visual data from cameras and depth sensors, natural language instructions from human operators, haptic feedback, and motion trajectories from manipulators. Integrating these varied data streams is essential for robots to build a comprehensive understanding of their environment and interact effectively. However, as highlighted by [jaeyong_deep_2015], manually designing features that bridge these modalities is extremely difficult due to differences in data formats and inherent complexities. For instance, visual data may be represented as images, point clouds, or voxel grids, each with unique dimensional structures, whereas language instructions are symbolic sequences that must be translated into a format suitable for robotic control. Similarly, manipulation trajectories can be represented as time series data of joint angles, end-effector positions, or force/torque measurements. This heterogeneity presents a substantial barrier to direct comparison or fusion of information across modalities.
Developing techniques that can automatically learn joint representations to capture relationships between these diverse modalities is crucial for advancing robot autonomy and intelligence. Additionally, as noted by [matthias_context_2018], context is essential for interpreting both linguistic and sensorimotor information. While embeddings are effective in capturing syntactic and some semantic relationships, they require contextualization to handle challenges such as word sense disambiguation and the integration of real-world knowledge, particularly in robotics applications.
@@ IV-B2 Methods for Multimodal Embeddings in Robotics
Deep neural networks are a powerful tool for learning shared embedding spaces that unify different modalities, mapping them into a common vector space where semantically related concepts are closer together. Various architectures are used for this purpose. For example, [yizhen_explainable_2021] discusses a two-stream architecture where separate networks process individual modalities, such as vision and language, before merging outputs into a joint representational space. Alternatively, a single network can concurrently process all modalities, directly learning a shared embedding space.
Training strategies for multimodal embeddings typically involve a two-stage process: an initial pre-training phase on large unimodal datasets, such as image-caption pairs or text corpora, to learn general features within each modality, followed by fine-tuning on smaller multimodal datasets to align representations across modalities. This approach proved effective in [shao-yen_multimodal_2019], where acoustic information was integrated into contextualized lexical embeddings to enhance emotion recognition through paralinguistic cues.
@@
Effective multimodal embedding training relies on carefully designed loss functions. Margin-based approaches, which aim to separate embeddings of dissimilar concepts while grouping similar ones, are commonly used. For instance, triplet loss, as discussed by [shahmohammadi_language_2022], teaches the model to position an anchor instance closer to a positive (semantically similar) instance than to a negative (semantically dissimilar) instance by a predefined margin. Similarly, contrastive loss, employed in [jaeyong_deep_2015], works by distinguishing between similar and dissimilar pairs, pushing dissimilar embeddings apart while pulling similar ones closer. These loss functions encourage a semantically structured embedding space that effectively captures relationships between modalities.
@@ IV-B3 Combining Point-Cloud Data, Natural Language Instructions, and Manipulation Trajectories
Embedding these diverse modalities into a shared space, as explored in [jaeyong_deep_2015], enables the establishment of meaningful connections across representations. For instance, robots can ground language instructions, such as ”grasp the blue object,” by linking the instruction to the corresponding point cloud of a blue object and its associated manipulation trajectory. This multimodal grounding enables generalization to novel objects and instructions, fostering more flexible and adaptive robotic behaviors.
@@ IV-B4 Applications of Multimodal Embeddings in Robotics
Multimodal embeddings enable robots to interact with and manipulate unfamiliar objects based on previous experiences. By learning a shared embedding space that includes object representations (e.g., point clouds), language instructions, and manipulation trajectories, as demonstrated by [jaeyong_deep_2015], robots can generalize their knowledge to novel objects. When encountering a new object alongside a language instruction, the robot can map the object’s representation into the shared embedding space, retrieving the most relevant manipulation trajectory aligned with the instruction. This capability allows robots to adapt to new situations and perform tasks without requiring explicit programming for every new object-action combination.
@@
Comprehending and executing complex tasks specified via natural language is a critical objective in robotics. Multimodal embeddings facilitate this by aligning language commands with corresponding action sequences. This enables robots to understand and perform multi-step tasks expressed in natural language, such as ”navigate to the table, pick up the book, and place it on the shelf.” With a multimodal embedding model, the robot can translate these instructions into a sequence of actions, including navigation, object recognition, grasping, and placement. This approach moves beyond simple, keyword-based command interpretation, fostering more flexible and sophisticated human-robot interactions. Future research, as suggested by [japa_question_2020], could enhance this capability by fine-tuning language models with attention mechanisms to improve the accuracy and robustness of mapping natural language instructions to specific actions.
@@
Evaluating multimodal embeddings in robotics requires a mix of real-world experiments and simulations. Real-world experiments provide a realistic measure of a model’s ability to generalize to new scenarios and handle the complexities of unstructured environments. However, these tests can be resource-intensive and time-consuming. Simulated environments offer a controlled and efficient platform to evaluate model performance across various tasks and conditions. For instance, [jaeyong_deep_2015] demonstrates how simulations can pre-train models before deploying them on physical robots. Common evaluation metrics include task success rate, execution time, and path efficiency.
@@
The field of multimodal embeddings for robotics presents ongoing challenges and exciting research directions. One primary challenge is scalability to high-dimensional sensor data. As robots incorporate advanced sensors like high-resolution cameras and lidar, the dimensionality of data increases significantly. Efficiently processing and embedding this high-dimensional data is essential for real-world applications. Another critical challenge is achieving real-time performance; real-time sensor data processing and action generation are vital for effective interaction in dynamic environments. Continued research is needed to create more efficient embedding models and architectures that support real-time robotic control. Exploring alternative embedding techniques, such as those discussed by [a_learning_2013], could yield scalable, efficient solutions for handling high-dimensional data in real-time applications.
@@ IV-C1 Motivation: Connecting Artificial and Biological Language Processing
The impressive capabilities of large language models (LLMs) in natural language processing tasks have prompted research into their internal representations and whether these mirror the way the human brain processes language. [yossi_analysis_2017] discusses the shift from word-level to sentence-level representations in language models, which reflects advances in understanding complex language representations. Exploring the connection between artificial and biological language processing is essential for developing AI that is more human-like and for gaining insights into the cognitive foundations of language comprehension. [richard_low-dimensional_2021] investigates how the structure of language representations correlates with human brain responses, highlighting the potential for identifying shared computational principles. Additionally, [matthias_context_2018] emphasizes the role of embeddings and contextual information, which could enhance our understanding of both artificial and biological language processing systems. This line of research not only enriches cognitive neuroscience but also informs the development of more robust and interpretable language models.